please help me with this age problem;

Harry is one year older than twice his sister’s age.
In three years, Harry will be nine years older than his sister. How old is Harry?

Answer by stanbon(75887)

Harry is one year older than twice his sister’s age.
In three years, Harry will be nine years older than his sister.
How old is Harry?
-----
Equations:
h = 2s + 1
h+3 = (s+3) + 9
---
Substitute for "h" and solve for "s":
2s+1 + 3 = s + 12
2s+4 = s+12
s = 8 (sister's age)
------
Solve for "h":
h = 2s+1
h = 2*8+1
h = 17 yrs (Harry's age)
=============================
Cheers,
Stan H.
=============================